Part A At what temperature (in。C) do hydrogen molecules have the same rms speed as nitrogen molecules at 140°C? Submit Request Answer ovide Feedback
Answer #1

The RMS velocity of a particle at temperature T and molecular mass M is given by

V​​​​​​rms​​​= √(3RT/M)

Therefore,

1) For Nitrogen,

M N= 14

T​​​​​​N = 140+273 = 413 K

Therefore, V​​​​​​rmsN =√(3R(413)/14) ------ { equation 1}

Now,

2) For Hydrogen,

M​​​​​​H = 2

T​​​​​​H=?

Therefore, V​​​​​​rmsH = √(3RT H/2) --------- { equation 2}

Hence, equating 1 and 2 -->

√(3R(413)/14)=√(3RT/2)

=> 413/14 = T/2

=> T H = 413/7 = 59 K = 59-273 °C = -214°C